OpenStudy (anonymous):

A circle with center (1,5) passes through the point (10,10). What is the area of the circle rounded to the nearest square unit?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

first solve for the radius of the circle

OpenStudy (anonymous):

I'm not sure how.

OpenStudy (anonymous):

wait ill draw it...

OpenStudy (anonymous):

Thanks but I don't know how to solve this at all.

OpenStudy (anonymous):

@mathmale

OpenStudy (wolf1728):

The radius is the hypotenuse of a triangle whose sides are 9 by 5 can you see that? can you solve the radius from that?

OpenStudy (anonymous):

plot the center and draw a line from the center up to the point where the circle pass through..

OpenStudy (anonymous):

I'm really quite lost here...

OpenStudy (mathmale):

Cubi-Cal: you wouldn't be in this course if you didn't already have at least some of the necessary background. The more you could review and share what you DO know that might apply to this problem, the better, and the more willing I would be to explain it from the point at which you feel stuck.

OpenStudy (wolf1728):

cubi cal - can you see how to calculate the radius?

OpenStudy (wolf1728):

see cubi-cal - we're here to help :-)

OpenStudy (anonymous):

can u atleast know the pythagorean theorem @Cubi-Cal ?

OpenStudy (mathmale):

There is a standard equation for a circle of radius r and center (h, k). Have you found and written down that equation for reference now and later?

OpenStudy (anonymous):

I'm sorry guys, I'm really locking up here. This is one my last problems for my final assignment so I'm kind of at the point where I'm brain dead. Yes! a^2+b^2=c^2

OpenStudy (wolf1728):

Okay cubi-cal - and c² would be the radius the other 2 sides are the 'x' length and the 'y' height

OpenStudy (mathmale):

Here's another viewpoint: If you have the center of a circle, EVERY point on that circle is the same distance from the center. Note the word "distance." Remember the distance formula? It's similar to your a^2 + b^2 = c^2, but must be modified if the center of the circle is not at (0,0).

OpenStudy (mathmale):

Drawing a circle could be very helpful, because then you could immediately see the line drawn from the center (1,5) to the point on the circle (10,10).

OpenStudy (anonymous):

Okay, hold on...

OpenStudy (anonymous):

Is the radius 9?

OpenStudy (mathmale):

Hold. I'd always prefer to see your work; otherwise I have to do it myself while you wait.

OpenStudy (mathmale):

Do you actually have that drawing in front of you?

OpenStudy (anonymous):

A very poor one, yes.

OpenStudy (anonymous):

9 is wrong....

OpenStudy (anonymous):

didi u try to use pythagorean theorem... the two sides are given by @wolf1728 .. which is 9 and 5

OpenStudy (mathmale):

We're going to try to find the radius. To do that, you're going to need to subtract one y-component from the other y-component. J-D got that by subtracting 5 from 10. Make certain you know why he did that. Next, you'll need to find the horizontal distance from x=1 to x=10. What is that distance?

OpenStudy (anonymous):

Oh dear. Um. 9?

OpenStudy (mathmale):

9, for what? Say it explicitly: the horiz. distance between the vertical lines x=1 and x=10 is 9. Yes. So now you have a right triangle with legs of 5 and 9, as J-D has already pointed out. Use the Pyth Thm to determine the length of the hypotenuse, which is also the radius of this circle.

OpenStudy (wolf1728):

So cubi-cal the radius² = 9² + 5² can you solve that?

OpenStudy (mathmale):

Yes: 9 squared plus 5 squared results in the SQUARE of the radius. What is the radius itself?

OpenStudy (mathmale):

You can find the radius itself if you wish. It just dawned on me that if you can find r^2 = a^2 + b^2, we can stop there. I'll explain why in a moment.

OpenStudy (anonymous):

106?

OpenStudy (mathmale):

For reference (please write this info down): The equation of a circle centered at (0,0) is x^2 + y^2 = r^2. The equation of a circle centered at (h, k) is \[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (mathmale):

Yes, great. The sqauare of the radius is indeed 106. So now you ahve everything you need tow rite th e equation of this circle: the center and the square of the radius. would you plese write this equatin now?

OpenStudy (wolf1728):

cubi-cal 106 is correct that is thhe radius squared

OpenStudy (anonymous):

How would I write the equation? Sorry.

OpenStudy (mathmale):

You've got to know the "standard equation of a circle centered at (h,k) with radius r:"\[(x-h)^2 + (y-k)^2 = r^2\] As before, r^2 = 106 and the center of the circle is given in the problem statement. Please write out the equation of this particular circle.

OpenStudy (mathmale):

If the center of this circle is at ( , ) (you fill in the blanks), then the x-coordinate of the center, h, has the value ( ). Likewise, the y-coord. is ( )

OpenStudy (mathmale):

Feel free to take the time necessary to go back to review what we've already discussed, above.

OpenStudy (anonymous):

(1,5) and (10,10) and (10,-10)

OpenStudy (mathmale):

The center of the circle is (1,5), meaning that the x-cooreinate, h, is 1. The x-coord. k, is .... What?

OpenStudy (mathmale):

x-coordinate (not cooreinate).

OpenStudy (mathmale):

Given that the x-coord. of the center is x = h = 1, the general equation of the circle becomes more specific:\[(x-1)^2 + (y-k)^2 = r^2\]...Note how I have replaced that 'h' by '1'.

OpenStudy (mathmale):

Finish writing this equation of your circle, please. k=?? r^2 = ??

OpenStudy (anonymous):

I'm feeling a bit overwhelmed. I'm so very sorry!

OpenStudy (wolf1728):

doing fine cubi - math male is doing very well too

OpenStudy (anonymous):

Thank you, you two are truly lifesavers!

OpenStudy (anonymous):

k though?

OpenStudy (wolf1728):

tnx

OpenStudy (anonymous):

I'm still REALLY unsure about it all.

OpenStudy (wolf1728):

on the phone at the moment - sorry :-(

OpenStudy (anonymous):

No worries! I'm gonna close this one and open a new one. I have two more questions then no more geometry!

OpenStudy (wolf1728):

I might be able to hepp you - depending on the phone

OpenStudy (anonymous):

Okay, no worries!

OpenStudy (anonymous):

WAIT I got it!!

OpenStudy (anonymous):

I just use distance formula!